### 11 Mar 2026 v1

# A note on Ramsey numbers for minors

Abstract
Let $R_{h}(k;\ell)$ be the smallest integer $n$ such that any edge coloring of a complete graph on $n$ vertices in $\ell$ colors results in a monochromatic graph with Hadwiger number $k$, i.e., a graph that could be transformed into a clique on $k$ vertices via a sequence of edge contractions and vertex deletions. We show that
$k\sqrt{\log k}(1+o(1))\leq R_{h}(k;2)\leq 1.031\cdot k\sqrt{\log k}(1+o(1))$
and for a constant $\beta=0.265656$...,
$R_{h}(k;\ell)=2\beta\ell k\sqrt{\log k}(1+o(1)).$
1 Introduction
The classical Ramsey number $R(k)$ is defined as the smallest integer $n$ such that any edge coloring of a complete graph $K_{n}$ on $n$ vertices with two colors contains a monochromatic $K_{k}$. This notion was generalised in many different ways, in particular “through parameter”. Specifically, when $\xi$ is some graph parameter, then $R_{\xi}(k)$ is defined as the smallest $n$ such that any edge coloring of $K_{n}$ in two colors contains a monochromatic subgraph $G$ with $\xi(G)=k$. So, $R(k)=R_{\omega}(k)$, where $\omega(G)$ is the clique number of $G$. One of the classical graph parameters is Hadwiger number $h(G)$ that is the largest $k$ so that $G$ contains a $K_{k}$-minor, that means that $K_{k}$ can be obtained from $G$ by edge-contractions and vertex-deletions. In this note, we shall discuss the Ramsey number for minors $R_{h}(k)$. Although it is very natural and complements extensive research on Hadwiger conjecture, it seems that it never was explicitly addressed in the literature. The Hadwiger conjecture [6] states that for any graph $G$, $h(G)\geq\chi(G)$, where $\chi(G)$ is the chromatic number of $G$, see for example Delcourt and Postle [5] and Hendrey, Norin, Steiner, and Turcotte [7]. A related density question is to bound $h(G)$ in terms of the average degree of $G$. For each $k\geq 2$, let
$c(k):={\rm min}\{c:|E(G)|\geq c|V(G)|\Rightarrow h(G)\geq k\}.$
The fact that one could take minimum and not just an infimum in the above definition is due to Mader [9]. Thomason [13] determined $c(k)$ asymptotically, see some related papers [1, 14, 15, 16]. Let $\lambda<1$ be the solution of the equation $1-\lambda+2\lambda\ln\lambda=0$ and and $\beta=(1-\lambda)/2\sqrt{\ln(1/\lambda)}/\sqrt{\log_{2}e}$. Here, $\beta=0.3190863.../\sqrt{\log_{2}e}=0.265656$....
Theorem 1.1 (Thomason [13]).
We have
$c(k)=\beta k\sqrt{\log_{2}k}(1+o(1)).$
The lower bound on $c(k)$ in the above theorem is achieved by the Erdős-Renyi random graph $G(n,p)$, whose Hadwiger number is well understood.
Theorem 1.2 (Bollobás, Catlin, and Erdős [4]).
For a fixed $p$ and q=1pq=1-p,
$h(G(n,p))=(1+o(1))n\left(\frac{\log(1/q)}{\log n}\right)^{1/2}.$
In addition, Thomason [13] proved that the Hadwider number for graphs whose connectivity is not too small is at least one of a random graph with the same density.
Theorem 1.3 (Theorem 4.1, Thomason [13]).
Let $0<\varepsilon<1$. There is $n_{0}=n_{0}(\varepsilon)$, such that if $G$ is a graph on $n$ vertices, $n>n_{0}$, of density $p$, and connectivity at least $n(\log\log\log n)/(\log\log n)$, then
$h(G)\geq\left\lceil(1-\varepsilon)n\left(\frac{\log(1/q^{*})}{\log n}\right)^{1/2}\right\rceil,$
where $q^{*}=\max\{1-p,(\log n)^{-1/\varepsilon}\}$.
Kostochka [8], Stiebitz [12], as well as Reed and Thomas [11] studied Hadwiger number for graphs and their complements in an “off-diagonal” setting when $h(G)$ is a fixed constant. For the rest on this note, the logarithms are base $2$, i.e., $\log=\log_{2}$ and $\beta$ is a constant defined above. We drop floors and ceilings where clear from context.
Let $R_{h}(k;\ell)$ be the smallest integer $n$ such that any edge coloring of $K_{n}$ in $\ell$ colors results in a monochromatic graph with Hadwiger number $k$. The following theorems bring known results together to get the bounds on $R_{h}(k;\ell)$. Recall that $R_{h}(k)=R_{h}(k;2)$.
Theorem 1.4.
We have
$k\sqrt{\log k}(1+o(1))\leq R_{h}(k)\leq 2\sqrt{\beta}k\sqrt{\log k}(1+o(1))\leq 1.031\cdot k\sqrt{\log k}(1+o(1)).$
Moreover, $R_{h}(3)=5$, $R_{h}(4)=7$, and for any $k\geq 5$, $2(k-2)\leq R_{h}(k)\leq 32(k-2)\lfloor\log(k-2)\rfloor+1.$
Theorem 1.5.
Let $\beta=0.2656568$... as above. For any $\ell\geq 1$, $R_{h}(3;\ell)=2\ell+1$. For any $\varepsilon>0$ there is $k_{0}=k_{0}(\varepsilon)$, such that for any $k\geq k_{0}$ there is $\ell_{0}=\ell_{0}(k,\varepsilon)$, so that for any $\ell>\ell_{0}$,
$2\beta\ell k\sqrt{\log k}(1-\epsilon)\leq R_{h}(k;\ell)\leq 2\beta\ell k\sqrt{\log k}(1+\epsilon).$
2 Proofs of the main results
First, we note that we can get an upper bound $R_{h}(k)\leq 4\beta k\sqrt{\log k}(1+o(1))\approx 1.0623k\sqrt{\log k}(1+o(1))$ from Theorem 1.1 immediately. Indeed, consider an arbitrary edge coloring of $K_{n}$ with two colors, where $n=4\beta k\sqrt{\log k}(1+o(1))$. Then the majority color class $G$ has size at least $\frac{n^{2}}{4}(1+o(1))=\beta k\sqrt{\log k}(1+o(1))n.$ From Theorem [13], we see that $h(G)\geq k$.
Now, we shall give a more refined argument to improve the constant term from $1.0623$... to $1.03108$.... Let $\varepsilon>0$ be a small constant and $n_{0}=n_{0}(\varepsilon)$ be sufficiently large, in particular satisfying assumptions of Theorem 1.3 and such that $(\log n)^{-1/\varepsilon}$ is sufficiently small, say at most $0.001$. Let $c=2\sqrt{\beta}+\varepsilon$. Note that for a small $\varepsilon$, $c>1/(1-2\varepsilon)$. Consider an arbitrary edge coloring of $K_{n}$ in red and blue, where $n=ck\sqrt{\log k}(1+o(1))$, and $k$ is sufficiently large so that $n>n_{0}$. Let without loss of generality red be a majority color class and $G$ be the graph whose edge set consists of all red edges. We define a function $x(n):=n(\log\log\log n)/(\log\log n)$.
Case 1. Graph $G$ has connectivity at least $x(n)$.
Let $p$ be the density of $G$. Note that $p$ and 1p1-p are larger than $(\log n)^{-1/\varepsilon}$, otherwise we can apply Theorem 1.1 to the red or the blue graph right away and get a monochromatic $K_{k}$-minor. Using the connectivity assumption and Theorem 1.3 with $q^{*}=(1-p)\leq 1/2$, we have $h(G)\geq(1-\varepsilon)n\sqrt{\log(1/2)/\log n}\geq ck(1-2\varepsilon)\geq k$.
Case 2. There is an induced subgraph $G^{\prime}$ of $G$ on at least $n^{\prime}=c^{\prime}n$ vertices, where $cc^{\prime}\geq 1/(1-3\varepsilon)$, minimum degree at least $x(n^{\prime})+k+1$, and density at least $1/2$.
If the connectivity of $G^{\prime}$ is at least $x(n^{\prime})$, then again by Theorem 1.3, we have that $h(G^{\prime})\geq(1-\varepsilon)c^{\prime}n/\sqrt{\log c^{\prime}n}\geq(1-2\varepsilon)cc^{\prime}k\sqrt{\log k}/\sqrt{\log k(1+o(1))}\geq k$. If the connectivity of $G^{\prime}$ is less than $x(n^{\prime})$, consider a smallest vertex cut $X$. By the condition on the minimum degree we see that any connected component of $G^{\prime}-X$ has size at least $k$. Thus there is a $K_{k,k}$ in the complement of $G$, i.e., a blue $K_{k}$-minor.
Consider a largest set $\{v_{1},\ldots,v_{y}\}$ of $y$ vertices such that $deg(v_{1})$ in $G$ and $deg(v_{i})$ in $G-\{v_{1},\ldots,v_{i-1}\}$ is at most $x(n)+k$, for any $2\leq i\leq y$. Let $G^{\prime}=G-\{v_{1},\ldots,v_{y}\}$. Then
$\frac{|E(G^{\prime})|}{|V(G^{\prime})|}\geq\frac{|E(G)|-y(x(n)+k)}{|V(G)|-y}\geq\frac{n/4-yx(n)/n-k/n}{1-y/n}.$
In particular, we have $|E(G^{\prime})|/|V(G^{\prime})|\geq|V(G^{\prime})|/4$ for any $y\geq 1$. If $y\geq n(1-c/(4\beta)+\varepsilon)$, then $|E(G^{\prime})|/|V(G^{\prime})|\geq(\beta+\varepsilon)k\sqrt{\log k}$. In this case by Theorem 1.1, $h(G)\geq k$. If $y<n(1-c/(4\beta)+\varepsilon)$, then $|V(G^{\prime})|>n(c/(4\beta)-\varepsilon)$. Thus, we see $G^{\prime}=G-\{v_{1},\ldots,v_{y}\}$ satisfies the conditions of Case 2 since $G^{\prime}$ has $c^{\prime}n$ vertices for $c^{\prime}\geq c/(4\beta)-\varepsilon$ and for sufficiently small $\varepsilon$ we have $cc^{\prime}\geq c^{2}/4\beta-c\varepsilon=1+\varepsilon/\sqrt{\beta}+\varepsilon^{2}/(4\beta)-2\sqrt{\beta}\varepsilon-\varepsilon^{2}>1/(1-3\varepsilon)$. This proves the upper bound.
For the lower bound we consider a random coloring of $K_{n}$, $n=k\sqrt{\log k}(1-\epsilon)$, where each edge is colored red with probability $1/2$ and blue with probability $1/2$. Each color class corresponds to $G(n,1/2)$. By Theorem 1.2, almost surely $h(G(n,1/2))<k$. This proves the first statement of the theorem.
Now we prove the bounds for small $k$ and give general, but not optimal bounds that hold for any $k$. To see the lower bound of 2(k2)2(k-2), consider a union of two red vertex-disjoint cliques on k2k-2 vertices each so that all other edges are blue. The upper bound follows from Theorem 1.16 by Bollobás [3], that claims that $c(k)\leq 8(k-2)\lfloor\log_{2}(k-2)\rfloor$. Indeed, let $G$ be a majority color class in a given red/blue coloring of the edges of $K_{n}$. It has at least n(n1)/4n(n-1)/4 edges. Since $(n-1)/4\geq 8(k-2)\lfloor\log_{2}(k-2)\rfloor$, we have that $h(G)\geq k$.
Finally, we treat the small cases $k\in\{3,4\}$. To see that $R_{h}(3)=5$, consider an arbitrary edge coloring of $K_{5}$ in two colors. A largest color class has at least $5$ edges and thus has a cycle, that is a $K_{3}$-minor. On the other hand, a coloring of $K_{4}$ where each color class is isomorphic to $P_{4}$ contains no cycles and thus no $K_{3}$-minors.
For $R_{h}(4)$, we shall use the fact that the largest number of edges in an $n$-vertex graph with no $K_{4}$-minor is at most 2n32n-3, for $n\geq 3$, see for example Bollobás [3]. To see that $R_{h}(4)>6$, consider a 6-cycle and three additional new edges forming a triangle, color this graph red and its complement blue. Since any $K_{4}$-minor has at least four vertices of degree at least three, we see that the color classes are $K_{4}$-minor free. We can immediately see that $R_{h}(4)\leq 8$. Indeed, in any 2-edge coloring of $K_{8}$, a majority color class has size at least $14>2\cdot 8-3$ and thus contains a $K_{4}$-minor. Now, we shall show a stronger bound $R_{h}(4)\leq 7$. Consider an arbitrary red/blue edge coloring of $K_{7}$ with no monochromatic $K_{4}$-minor and let $G$ be a graph whose edge set consists of the red edges. We make several observations. As above $10\leq|E(G)|\leq 11$. Both $G$ and $\overline{G}$ are not bipartite since otherwise one of the parts has size at least $4$. By a special case of Hadwiger conjecture, $\chi(G)=\chi(\overline{G})=3$. Thus $G$ and $\overline{G}$ are three-partite graphs with parts of sizes at most three, i.e., of sizes $3,3,1$ or $3,2,2$. There is no monochromatic $K_{3,3}$ without an edge. If one color class, say blue, has two vertex disjoint triangles, then there are two blue edges between them or three edges forming a $P_{4}$ between them. By considering the remaining vertex, one can see that there is a monochromatic copy of $K_{4}$-minor, a contradiction. Both the red and the blue graphs are three-chromatic with color classes of sizes $3,2,2$. This leaves a final not so long case analysis to demonstrate the existence of a monochromatic $K_{4}$-minor, a contradiction. ∎
To see that $R_{h}(3;\ell)\geq 2\ell+1$, consider a decomposition of $K_{2\ell}$ into $\ell$ Hamiltonian paths. Since any $K_{3}$-minor contains a cycle, we have that this coloring has no monochromatic $K_{3}$-minors. To see that $R_{h}(3;\ell)\leq 2\ell+1$, consider an arbitrary edge-coloring of $K_{2\ell+1}$ in $\ell$ colors. Then some color class has at least $2\ell+1$ edges and thus contains a cycle, that is a $K_{3}$-minor.
The upper bound of the theorem holds in a stronger form, for any $\ell$. Let $n=R_{h}(k;\ell)-1$ and consider an edge-coloring of $K_{n}$ in $\ell$ colors with no monochromatic $K_{k}$-minor. From Theorem 1.1, we know that any $K_{k}$-minor free graph on $n$ vertices has at most $\beta k\sqrt{\log k}n(1+o_{k}(1))$ edges. Since the total number of edges on $n$ vertices is $\binom{n}{2}$, we have that $\binom{n}{2}\leq\ell\cdot\beta k\sqrt{\log k}n(1+o_{k}(1))$. This implies that $n\leq\ell\cdot 2\beta k\sqrt{\log k}(1+o_{k}(1))$.
For the lower bound, let $\varepsilon$ be a small constant, $0<\varepsilon<1$, $k$ be sufficiently large. Consider a graph $H^{\prime}$ on $k\sqrt{\log k}$ vertices and $m=\beta k^{2}\log k(1-\varepsilon^{2}/5)$ edges that is $K_{k}$-minor free, that exists by Theorem 1.1. Let $H$ be obtained from $H^{\prime}$ by attaching a pendant edge. Wilson [17] showed that $K_{n}$ is a pairwise edge-disjoint union of copies of $H$ if $n$ is sufficiently large, $|E(H)|$ divides $\binom{n}{2}$, and the greatest common divisor of the degrees of $H$ divides n1n-1. Since there is a vertex of degree $1$ in $H$, the last condition is satisfied for any $n\geq 2$. Let $\ell$ be sufficiently large, so that for some $n$ satisfying $\ell\cdot 2\beta k\sqrt{\log k}(1-\varepsilon)\leq n\leq\ell\cdot 2\beta k\sqrt{\log k}(1-\varepsilon/2)$, $n$ satisfies conditions of Wilson’s theorem, in particulst $|E(H)|$ divides $\binom{n}{2}$. Then $H$ decomposes $K_{n}$. Alon and Yuster [2] proved that if a fixed graph $H$ decomposes $K_{n}$, then there is a coloring of $E(K_{n})$ in at most (1+o(1))(n1)|V(H)|/(2|E(H)|)(1+o(1))(n-1)|V(H)|/(2|E(H)|) colors, such that each color class is a pairwise vertex-disjoint union of copies of $H$. So, we take $\ell$ sufficiently large that $o(1)$ term is at most $\varepsilon/2$. That is, every color class is almost an $H$-factor and it is $K_{k}$-minor free. Since the number of colors is at most $(1+\varepsilon/2)(n-1)|V(H)|/(2|E(H)|)\leq(1+\varepsilon/2)\ell\cdot 2\beta k\sqrt{\log k}(1-\varepsilon/2)/(2\beta k\sqrt{\log k}(1-\varepsilon^{2}/5))\leq\ell$, we have a coloring of $E(K_{n})$ in $\ell$ colors with no monochromatic copy of a $K_{k}$-minor. ∎
3 Concluding remarks
We showed that $R_{h}(k;2)=ck\sqrt{\log k}(1+o(1))$, where $c$ is a constant, $1\leq c\leq 1.03$.... We believe that $c=1$. It would be interesting to find an easy direct proof of an upper bound on $R_{h}(k;2)\leq ck\sqrt{\log k}$, for some, maybe large constant $c$, but avoiding the density Theorem 1.1.
To prove the lower bound on $R_{h}(k;\ell)$, one can also use the following packing result by taking $G$ being an $H$-factor, where $H$ is a graph on $k\sqrt{\log k}(1+o(1))$ vertices, density $\beta k\sqrt{\log k}$ and no $K_{k}$-minor. Pack $K_{n(1-\varepsilon)}$ with copies of $G$, give each such copy of $G$ own color. Decompose the remaining edges of $K_{n}$ into forests using Nash-Williams theorem and give each such forest a new own color.
Theorem 3.1 (Messuti, Rödl, Schacht [10]).
For any $\varepsilon>0$ and any $\Delta\in\mathbb{N}$, for any non-trivial minor closed family $\mathcal{G}$ there is $n_{0}$ such that for any $n\geq n_{0}$ and any integer $\ell$ the following holds. If $G\in\mathcal{G}$, $\Delta(G)\leq\Delta$ and $|V(G)|\leq n$, and $\ell|E(G)|\leq\binom{n}{2}$ then $\ell$ copies of $G$ pack in $K_{(1+\varepsilon)n}$.
4 Acknowledgements
The author is grateful to Noga Alon for several very fruitful discussions and suggestions. She also thanks Jan Malte Weirich for discussions.

### 23 Mar 2026 v2

@@
 Abstract
-Let $R_{h}(k;\ell)$ be the smallest integer $n$ such that any edge coloring of a complete graph on $n$ vertices in $\ell$ colors results in a monochromatic graph with Hadwiger number $k$, i.e., a graph that could be transformed into a clique on $k$ vertices via a sequence of edge contractions and vertex deletions.
-We show that
-$k\sqrt{\log k}(1+o(1))\leq R_{h}(k;2)\leq 1.031\cdot k\sqrt{\log k}(1+o(1))$
-and for a constant $\beta=0.265656$...,
-$R_{h}(k;\ell)=2\beta\ell k\sqrt{\log k}(1+o(1)).$
+Let $R_{h}(k;\ell)$ be the smallest integer $n$ such that any edge coloring of a complete graph on $n$ vertices in $\ell$ colors results in a monochromatic $K_{k}$-minor, in other words, a graph with Hadwiger number $k$, i.e., a graph that could be transformed into a clique $K_{k}$ on $k$ vertices via a sequence of edge contractions and vertex deletions.
+More generally, for a graph $F$ and integer $\ell$ let $R_{h}(F;\ell)$ be the smallest integer $n$ such that any edge coloring of a complete graph on $n$ vertices in $\ell$ colors results in a monochromatic $F$-minor.
+In 2001 Thomason and in 2005 Myers and Thomason asymptotically determined the extremal numbers for clique minors and $F$-minors, respectively.
+They found the respective explicitly computable leading constants $\beta=0.265656$...
+and $\gamma(F)\cdot\beta$ for these extremal numbers.
+We determine $R_{h}(F;2)$ for every graph $F$ as
+$R_{h}(F;2)=(\gamma(F)+o(1))|V(F)|\sqrt{\log_{2}(|V(F)|)},$
+where the $o(1)$-term tends to zero as $|V(F)|\rightarrow\infty$. In particular,
+$R_{h}(k;2)=(1+o(1))k\sqrt{\log_{2}k}.$
+When $\ell\gg k\gg 1$, we show that
+$R_{h}(k;\ell)=(2\beta+o(1))\ell k\sqrt{\log_{2}k}.$
 1 Introduction
@@
 So, $R(k)=R_{\omega}(k)$, where $\omega(G)$ is the clique number of $G$.
-One of the classical graph parameters is Hadwiger number $h(G)$ that is the largest $k$ so that $G$ contains a $K_{k}$-minor, that means that $K_{k}$ can be obtained from $G$ by edge-contractions and vertex-deletions.
+One of the classical graph parameters is the Hadwiger number $h(G)$ which is the largest $k$ so that $G$ contains a $K_{k}$-minor, that means that $K_{k}$ can be obtained from $G$ by edge-contractions and vertex-deletions.
 In this note, we shall discuss the Ramsey number for minors $R_{h}(k)$.
-Although it is very natural and complements extensive research on Hadwiger conjecture, it seems that it never was explicitly addressed in the literature.
-The Hadwiger conjecture [6] states that for any graph $G$, $h(G)\geq\chi(G)$, where $\chi(G)$ is the chromatic number of $G$, see for example Delcourt and Postle [5] and Hendrey, Norin, Steiner, and Turcotte [7].
+Although it is very natural and complements extensive research on Hadwiger’s conjecture, it seems that it never was explicitly addressed in the literature.
+The Hadwiger conjecture [7] states that for any graph $G$, $h(G)\geq\chi(G)$, where $\chi(G)$ is the chromatic number of $G$, see for example Delcourt and Postle [5], see also [13] for a survey.
 A related density question is to bound $h(G)$ in terms of the average degree of $G$.
@@
 We have
-$c(k)=\beta k\sqrt{\log_{2}k}(1+o(1)).$
+$c(k)=(1+o(1))\beta k\sqrt{\log_{2}k}.$
 The lower bound on $c(k)$ in the above theorem is achieved by the Erdős-Renyi random graph $G(n,p)$, whose Hadwiger number is well understood.
@@
 $h(G(n,p))=(1+o(1))n\left(\frac{\log(1/q)}{\log n}\right)^{1/2}.$
-In addition, Thomason [13] proved that the Hadwider number for graphs whose connectivity is not too small is at least one of a random graph with the same density.
+In addition, Thomason [15] proved that the Hadwiger number for graphs whose connectivity is not too small is at least one of a random graph with the same density.
+Here, the density of a graph $G$ is $|E(G)|/\binom{|V(G)|}{2}$.
 Theorem 1.3 (Theorem 4.1, Thomason [15]).
@@
 where $q^{*}=\max\{1-p,(\log n)^{-1/\varepsilon}\}$.
-Kostochka [8], Stiebitz [12], as well as Reed and Thomas [11] studied Hadwiger number for graphs and their complements in an “off-diagonal” setting when $h(G)$ is a fixed constant.
-For the rest on this note, the logarithms are base $2$, i.e., $\log=\log_{2}$ and $\beta$ is a constant defined above.
+Each of Theorems 1.1, 1.2, 1.3 have been generalized to arbitrary minors $F$ by Myers and Thomason [11], who in particular determined $c(F)$, where
+$c(F):=\inf\{c:|E(G)|\geq c|V(G)|\Rightarrow G\mbox{ has an }\mbox{$F$-minor}\}.$
+To state these more general results, we need to introduce a parameter $\gamma(F)\in[0,1]$ associated with any graph which was defined by Myers and Thomason [11] as the optimal value of an optimization problem over vertex-weightings as follows.
+Definition 1.4.
+Let $F$ be a graph on $k$ vertices. Then we define
+$\gamma(F)=\min\left\{\frac{1}{k}\sum_{v\in V(F)}w(v)\bigg|w:V(F)\rightarrow[0,\infty)\text{ and }\sum_{uv\in E(F)}k^{-w(u)w(v)}\leq k\right\}.$
+Myers and Thomason [11] determined $\gamma(H)$ precisely for several interesting graphs.
+In particular, they showed that $\gamma(K_{k})=1$ for every $k$, gave an explicit formula for $\gamma(F)$ for every complete multi-partite graph $F$, showed that every graph $F$ on $k$ vertices with average degree $k^{\tau}$ satisfies $\gamma(F)\leq\sqrt{\tau}$, and most interestingly that almost all graphs $F$ on $k$ vertices with this average degree satisfy $\gamma(F)=\sqrt{\tau}-o(1)$.
+With this notation at hand, we can now state the aforementioned generalizations of Theorems 1.1, 1.2, 1.3 to arbitrary minors.
+Theorem 1.5 (Theorem 2.2, Myers and Thomason [11]).
+Let $F$ be a graph on $k$ vertices. Then
+$c(F)=(\gamma(F)\beta+o(1))k\sqrt{\log_{2}k},$
+where the o(1)-term tends to 0 as $k\rightarrow\infty$.
+Theorem 1.6 (Theorem 2.3, Myers and Thomason[11]).
+For every $\varepsilon>0$ there exists $k_{0}=k_{0}(\varepsilon)$ such that for every $k\geq k_{0}$, every graph $F$ with $k$ vertices satisfying $\gamma(F)\geq\varepsilon$, every $p\in[\varepsilon,1-\varepsilon]$, and for every number $n\leq\gamma(F)k\sqrt{\log_{1/(1-p)}(k)}$, we have that $F$ is a minor of $G(n,p-\varepsilon)$ with probability less than $\varepsilon$.
+Theorem 1.7 (Theorem 2.5, Myers and Thomason [11]).
+For every $\varepsilon>0$ there exists $k_{0}=k_{0}(\varepsilon)$ such that for every $k\geq k_{0}$, every graph $F$ with $k$ vertices satisfying $\gamma(F)\geq\varepsilon$, every $p\in[\varepsilon,1-\varepsilon]$, and for every number $n\geq\gamma(F)k\sqrt{\log_{1/(1-p)}(k)}$, the following holds.
+If $G$ is a graph on $n$ vertices, with density $p+\varepsilon$ and connectivity at least $n(\log\log\log n)/(\log\log n)$, then $F$ is a minor of $G$.
+Kostochka [8], Stiebitz [14], as well as Reed and Thomas [12] studied the Hadwiger number for graphs and their complements in an “off-diagonal” setting when $h(G)$ is a fixed constant.
+For the rest of this note, the logarithms are base $2$, i.e., $\log=\log_{2}$ and $\beta$ is a constant defined above.
 We drop floors and ceilings where clear from context.
-Let $R_{h}(k;\ell)$ be the smallest integer $n$ such that any edge coloring of $K_{n}$ in $\ell$ colors results in a monochromatic graph with Hadwiger number $k$.
-The following theorems bring known results together to get the bounds on $R_{h}(k;\ell)$.
-Recall that $R_{h}(k)=R_{h}(k;2)$.
+Let $R_{h}(k;\ell)$ be the smallest integer $n$ such that any edge coloring of $K_{n}$ in $\ell$ colors results in a monochromatic graph with Hadwiger number $k$ (we also abbreviate $R_{h}(k):=R_{h}(k;2)$).
+More generally, for a graph $F$ let us denote by $R_{h}(F;\ell)$ the smallest $n$ such that every $\ell$-edge coloring of $K_{n}$ results in a monochromatic $F$-minor.
+The following theorems leverage the previously stated results of Myers and Thomason to asymptotically determine $R_{h}(F;2)$ up to lower-order terms for all graphs $F$ and in particular $R_{h}(k)$, and also to asymptotically determine $R_{h}(k;\ell)$ when $\ell$ is much larger than $k$.
 Theorem 1.8.
-We have
-$k\sqrt{\log k}(1+o(1))\leq R_{h}(k)\leq 2\sqrt{\beta}k\sqrt{\log k}(1+o(1))\leq 1.031\cdot k\sqrt{\log k}(1+o(1)).$
-Moreover, $R_{h}(3)=5$, $R_{h}(4)=7$, and for any $k\geq 5$, $2(k-2)\leq R_{h}(k)\leq 32(k-2)\lfloor\log(k-2)\rfloor+1.$
+For every graph $F$ on $k$ vertices we have
+$R_{h}(F;2)=(\gamma(F)+o(1))k\sqrt{\log k}$
+where the $o(1)$-term tends to zero as $k\rightarrow\infty$. In particular, we have
+$R_{h}(k)=(1+o(1))k\sqrt{\log k}.$
+Moreover, $R_{h}(3)=5$, $R_{h}(4)=7$, and for any $k\geq 5$, $2k-2\leq R_{h}(k)\leq 32(k-2)\lfloor\log(k-2)\rfloor+1.$
 Theorem 1.9.
@@
 For any $\varepsilon>0$ there is $k_{0}=k_{0}(\varepsilon)$, such that for any $k\geq k_{0}$ there is $\ell_{0}=\ell_{0}(k,\varepsilon)$, so that for any $\ell>\ell_{0}$,
-$2\beta\ell k\sqrt{\log k}(1-\epsilon)\leq R_{h}(k;\ell)\leq 2\beta\ell k\sqrt{\log k}(1+\epsilon).$
+$2\beta\ell k\sqrt{\log k}(1-\varepsilon)\leq R_{h}(k;\ell)\leq 2\beta\ell k\sqrt{\log k}(1+\varepsilon).$
 2 Proofs of the main results
-First, we note that we can get an upper bound $R_{h}(k)\leq 4\beta k\sqrt{\log k}(1+o(1))\approx 1.0623k\sqrt{\log k}(1+o(1))$ from Theorem 1.1 immediately.
-Indeed, consider an arbitrary edge coloring of $K_{n}$ with two colors, where $n=4\beta k\sqrt{\log k}(1+o(1))$.
-Then the majority color class $G$ has size at least $\frac{n^{2}}{4}(1+o(1))=\beta k\sqrt{\log k}(1+o(1))n.$ From Theorem [13], we see that $h(G)\geq k$.
-Now, we shall give a more refined argument to improve the constant term from $1.0623$...
-to $1.03108$....
-Let $\varepsilon>0$ be a small constant and $n_{0}=n_{0}(\varepsilon)$ be sufficiently large, in particular satisfying assumptions of Theorem 1.3 and such that $(\log n)^{-1/\varepsilon}$ is sufficiently small, say at most $0.001$.
-Let $c=2\sqrt{\beta}+\varepsilon$.
-Note that for a small $\varepsilon$, $c>1/(1-2\varepsilon)$.
-Consider an arbitrary edge coloring of $K_{n}$ in red and blue, where $n=ck\sqrt{\log k}(1+o(1))$, and $k$ is sufficiently large so that $n>n_{0}$.
-Let without loss of generality red be a majority color class and $G$ be the graph whose edge set consists of all red edges.
-We define a function $x(n):=n(\log\log\log n)/(\log\log n)$.
-Case 1. Graph $G$ has connectivity at least $x(n)$.
-Let $p$ be the density of $G$.
-Note that $p$ and 1p1-p are larger than $(\log n)^{-1/\varepsilon}$, otherwise we can apply Theorem 1.1 to the red or the blue graph right away and get a monochromatic $K_{k}$-minor.
-Using the connectivity assumption and Theorem 1.3 with $q^{*}=(1-p)\leq 1/2$, we have $h(G)\geq(1-\varepsilon)n\sqrt{\log(1/2)/\log n}\geq ck(1-2\varepsilon)\geq k$.
-Case 2.
-There is an induced subgraph $G^{\prime}$ of $G$ on at least $n^{\prime}=c^{\prime}n$ vertices, where $cc^{\prime}\geq 1/(1-3\varepsilon)$, minimum degree at least $x(n^{\prime})+k+1$, and density at least $1/2$.
-If the connectivity of $G^{\prime}$ is at least $x(n^{\prime})$, then again by Theorem 1.3, we have that $h(G^{\prime})\geq(1-\varepsilon)c^{\prime}n/\sqrt{\log c^{\prime}n}\geq(1-2\varepsilon)cc^{\prime}k\sqrt{\log k}/\sqrt{\log k(1+o(1))}\geq k$.
-If the connectivity of $G^{\prime}$ is less than $x(n^{\prime})$, consider a smallest vertex cut $X$.
-By the condition on the minimum degree we see that any connected component of $G^{\prime}-X$ has size at least $k$.
-Thus there is a $K_{k,k}$ in the complement of $G$, i.e., a blue $K_{k}$-minor.
-Consider a largest set $\{v_{1},\ldots,v_{y}\}$ of $y$ vertices such that $deg(v_{1})$ in $G$ and $deg(v_{i})$ in $G-\{v_{1},\ldots,v_{i-1}\}$ is at most $x(n)+k$, for any $2\leq i\leq y$.
-Let $G^{\prime}=G-\{v_{1},\ldots,v_{y}\}$.
-Then
-$\frac{|E(G^{\prime})|}{|V(G^{\prime})|}\geq\frac{|E(G)|-y(x(n)+k)}{|V(G)|-y}\geq\frac{n/4-yx(n)/n-k/n}{1-y/n}.$
-In particular, we have $|E(G^{\prime})|/|V(G^{\prime})|\geq|V(G^{\prime})|/4$ for any $y\geq 1$.
-If $y\geq n(1-c/(4\beta)+\varepsilon)$, then $|E(G^{\prime})|/|V(G^{\prime})|\geq(\beta+\varepsilon)k\sqrt{\log k}$.
-In this case by Theorem 1.1, $h(G)\geq k$.
-If $y<n(1-c/(4\beta)+\varepsilon)$, then $|V(G^{\prime})|>n(c/(4\beta)-\varepsilon)$.
-Thus, we see $G^{\prime}=G-\{v_{1},\ldots,v_{y}\}$ satisfies the conditions of Case 2 since $G^{\prime}$ has $c^{\prime}n$ vertices for $c^{\prime}\geq c/(4\beta)-\varepsilon$ and for sufficiently small $\varepsilon$ we have $cc^{\prime}\geq c^{2}/4\beta-c\varepsilon=1+\varepsilon/\sqrt{\beta}+\varepsilon^{2}/(4\beta)-2\sqrt{\beta}\varepsilon-\varepsilon^{2}>1/(1-3\varepsilon)$.
-This proves the upper bound.
-For the lower bound we consider a random coloring of $K_{n}$, $n=k\sqrt{\log k}(1-\epsilon)$, where each edge is colored red with probability $1/2$ and blue with probability $1/2$.
+We start by proving the first part of the theorem concerning the asymptotic behavior of $R_{h}(F;2)$, which then also implies the asymptotic behavior of $R_{h}(k)$, since $R_{h}(k)=R_{h}(K_{k};2)$ and $\gamma(K_{k})=1$.
+Formally, we need to prove that for every $\varepsilon>0$ and every sufficiently large number $k$ (in terms of $\varepsilon$), every graph $F$ on $k$ vertices satisfies (denoting $\gamma=\gamma(F)$ in the following):
+$(\gamma-\varepsilon)k\sqrt{\log_{2}k}\leq R(F;2)\leq(\gamma+\varepsilon)k\sqrt{\log_{2}k}.$
+Suppose first that $\gamma\leq\varepsilon/2$.
+Then the claimed lower bound holds trivially for every $k$.
+To see that the upper bound holds for every sufficiently large $k$ in this case, consider a coloring of the edges of $K_{n}$ in two colors for $n\geq\lfloor(\gamma+\varepsilon)k\sqrt{\log_{2}k}\rfloor$.
+A larger of the two color classes corresponds to a graph $G$ satisfying $|E(G)|/|V(G)|\geq(n(n-1)/4)/n=(n-1)/4\geq(\beta\gamma+\varepsilon/10)k\sqrt{\log_{2}k}$ for every large enough $k$, where the last inequality holds since $\gamma+\varepsilon\geq 4(\beta\gamma+\varepsilon/8)$.
+Hence, by Theorem 1.5, $G$ contains $F$ as a minor given that $k$ is sufficiently large, as desired.
+Moving on, we may thus assume that $\gamma\geq\varepsilon/2$. Furthermore, we may trivially assume w.l.o.g. that $\varepsilon<1/2$.
+We first prove the lower bound on $R_{h}(F;2)$.
+To do so, we consider a random coloring of $K_{n}$, where $n=\lfloor(\gamma-\varepsilon)k\sqrt{\log k}\rfloor$, and where each edge is independently colored red with probability $1/2$ and blue with probability $1/2$.
 Each color class corresponds to $G(n,1/2)$.
-By Theorem 1.2, almost surely $h(G(n,1/2))<k$.
-This proves the first statement of the theorem.
+Choose $\varepsilon^{\ast}>0$ (only dependent on $\varepsilon)$ small enough such that $\varepsilon^{\ast}<\varepsilon/2$, $1/2+\varepsilon^{\ast}\leq 1-\varepsilon$, and $1/\sqrt{\log_{2}(1/(1/2-\varepsilon^{\ast}))}\geq 1-\varepsilon$.
+Now, let us set $p^{\ast}:=1/2+\varepsilon^{\ast}$ and note that by our choice of $\varepsilon^{\ast}$, we have $\gamma\geq\varepsilon/2\geq\varepsilon^{\ast}$, $p^{\ast}\in[\varepsilon^{\ast},1-\varepsilon^{\ast}]$ and
+$n\leq(\gamma-\varepsilon)k\sqrt{\log_{2}k}\leq(1-\varepsilon)\gamma k\sqrt{\log_{2}k}\leq\frac{\gamma k\sqrt{\log_{2}k}}{\sqrt{\log_{2}(1/(1/2-\varepsilon^{\ast}))}}=\gamma k\sqrt{\log_{1/(1-p^{\ast})}k}.$
+Hence, the preconditions of Theorem 1.6 with parameters $\varepsilon^{\ast},p^{\ast}$ are met, and we find for every sufficiently large $k$ that $G(n,1/2)=G(n,p^{\ast}-\varepsilon^{\ast})$ contains $F$ as a minor with probability less than $\varepsilon^{\ast}<1/2$.
+Hence, with positive probability, both color classes of our random edge-coloring of $K_{n}$ do not contain $F$ as a minor, showing that $R_{h}(F;2)\geq n+1>(\gamma-\varepsilon)k\sqrt{\log_{2}k}$.
+This concludes the proof of the desired lower bound.
+Next, let us prove the upper bound.
+So let $k$ be sufficiently large in terms of $\varepsilon$ (to be specified later) and let a coloring of the edges of $K_{n}$ with colors red and blue be given to us, where $n=\lfloor(\gamma(F)+\varepsilon)k\sqrt{\log_{2}k}\rfloor$.
+Let $G$ denote the graph on the same vertex set as $K_{n}$ formed by the edges of a color class with the most edges, w.l.o.g.
+let us say that this color class consists of the red edges.
+Let $\varepsilon^{\prime}>0$ be chosen sufficiently small, only depending on $\varepsilon$, such that $\varepsilon^{\prime}<\min\{\varepsilon/8,1/100\}$ and such that $\frac{1+10\varepsilon^{\prime}}{\log_{2}\left(2/(1+10\varepsilon^{\prime})\right)}<(1+\varepsilon/2)^{2}.$
+Let now $H$ denote the $\lceil\varepsilon^{\prime}n\rceil$-core of $G$, i.e., the (unique) maximal induced subgraph of $G$ with minimum degree at least $\varepsilon^{\prime}n$.
+To see that $H$ is well-defined, recall that for every $d\in\mathbb{R}^{+}$, every graph of average degree at least $d$ contains a non-empty subgraph of minimum degree at least $d/2$.
+Since $G$ contains at least half of the edges of $K_{n}$, it has average degree at least (n1)/2(n-1)/2 and hence contains a subgraph with minimum degree at least $(n-1)/4>\varepsilon^{\prime}n$ for $k$ sufficiently large.
+Note that $H$ can be obtained from $G$ by repeatedly deleting vertices of degree less than $\varepsilon^{\prime}n$.
+Hence, we have $|E(H)|\geq|E(G)|-\varepsilon^{\prime}n^{2}\geq n(n-1)/4-\varepsilon^{\prime}n^{2}\geq(1/4-2\varepsilon^{\prime})n^{2}$ for $k$ large enough.
+This in particular implies that $x:=|V(H)|\geq\sqrt{2|E(H)|}\geq\sqrt{1/2-4\varepsilon^{\prime}}\cdot n$.
+In the following, let $\alpha:=x/n\in\left[\sqrt{1/2-4\varepsilon^{\prime}},1\right]$ and let $p:=|E(H)|/\binom{x}{2}$ denote the density of $H$.
+Note that from the above, we immediately obtain that
+$p\geq(1/4-2\varepsilon^{\prime})n^{2}/\binom{x}{2}\geq(1/2-4\varepsilon^{\prime})(n/x)^{2}=1/(2\alpha^{2})-(4\varepsilon^{\prime})/\alpha^{2}\geq 1/2-4\varepsilon^{\prime}.$
+Moving on, we now distinguish two cases.
+Case 1. $H$ has vertex-connectivity at least $n(\log\log\log n)/(\log\log n)$.
+So, $H$ has $x$ vertices, $x\geq\sqrt{1/2-4\varepsilon^{\prime}}\cdot n$, it has connectivity at least $x(\log\log\log x)/(\log\log x)$, and it has density $p\geq 1/2-4\varepsilon^{\prime}\geq 2\varepsilon^{\prime}$ (by our choice of $\varepsilon^{\prime}$).
+We shall show that $H$ satisfies the conditions of Theorem 1.7 and thus has an $F$-minor.
+Let $p^{\prime}:=p-\varepsilon^{\prime}$. Then $p^{\prime}\in[\varepsilon^{\prime},1-\varepsilon^{\prime}]$. Also, recall that $\gamma\geq\varepsilon/2\geq\varepsilon^{\prime}$.
+Next, we claim that $x\geq\gamma k\sqrt{\log_{1/(1-p^{\prime})}(k)}$ for large enough $k$. Indeed, we have
+$x=\alpha n\geq\alpha\left(\gamma+\frac{\varepsilon}{2}\right)k\sqrt{\log_{2}(k)}\geq\alpha\gamma\left(1+\frac{\varepsilon}{2}\right)k\sqrt{\log_{2}(k)}=\frac{(1+\frac{\varepsilon}{2})\alpha}{\sqrt{\log_{(1/(1-p^{\prime}))}2}}\cdot\gamma k\sqrt{\log_{1/(1-p^{\prime})}(k)}.$
+It remains to show that $(1+\varepsilon/2)\alpha\sqrt{\log_{2}(1/(1-p^{\prime}))}\geq 1$.
+To do so, we start by recalling that $p^{\prime}=p-\varepsilon^{\prime}\geq\frac{1}{2\alpha^{2}}-\frac{4\varepsilon^{\prime}}{\alpha^{2}}-\varepsilon^{\prime}$.
+Since furthermore $\alpha\geq\sqrt{\frac{1}{2}-4\varepsilon^{\prime}}$ and $\varepsilon^{\prime}<1/100$, we find that in fact $p^{\prime}\geq\frac{1}{2\alpha^{2}}-\frac{4\varepsilon^{\prime}}{\frac{1}{2}-4\varepsilon^{\prime}}-\varepsilon^{\prime}\geq\frac{1}{2\alpha^{2}}-10\varepsilon^{\prime}$.
+Rearranging yields $\alpha\geq\frac{1}{\sqrt{2p^{\prime}+20\varepsilon^{\prime}}}$, and so we obtain
+$(1+\varepsilon/2)\alpha\sqrt{\log_{2}(1/(1-p^{\prime}))}\geq(1+\varepsilon/2)\sqrt{\frac{\log_{2}(1/(1-p^{\prime}))}{2p^{\prime}+20\varepsilon^{\prime}}}.$
+Since $p^{\prime}\geq 1/2-5\varepsilon^{\prime}$, we furthermore have
+$2p^{\prime}+20\varepsilon^{\prime}\leq 2p^{\prime}\left(1+\frac{10\varepsilon^{\prime}}{1/2-5\varepsilon^{\prime}}\right)=2p^{\prime}\frac{1+10\varepsilon^{\prime}}{1-10\varepsilon^{\prime}}.$
+Plugging this in, it follows that
+$(1+\varepsilon/2)\alpha\sqrt{\log_{2}(1/(1-p^{\prime}))}\geq(1+\varepsilon/2)\sqrt{\frac{\log_{2}(1/(1-p^{\prime}))}{2p^{\prime}}}\cdot\sqrt{\frac{1-10\varepsilon^{\prime}}{1+10\varepsilon^{\prime}}}.$
+It is not difficult to check that the real-valued function $f:(0,1)\rightarrow\mathbb{R}_{+}$, $f(y):=\frac{\log_{2}(1/(1-y))}{2y}$ is monotonically increasing in $y$.
+In particular, $f(p^{\prime})\geq f(1/2-5\varepsilon^{\prime})$, yielding
+$(1+\varepsilon/2)\alpha\sqrt{\log_{2}(1/(1-p^{\prime}))}\geq(1+\varepsilon/2)\sqrt{\frac{\log_{2}\left(2/(1+10\varepsilon^{\prime})\right)}{1+10\varepsilon^{\prime}}}\geq 1,$
+where we used our choice of $\varepsilon^{\prime}$ in the last step.
+This is the desired inequality, and so we have shown that indeed, $H$ has $x\geq\gamma k\sqrt{\log_{1/(1-p^{\prime})}(k)}$ vertices for large enough $k$.
+Note that $H$ has density $p=p^{\prime}+\varepsilon^{\prime}$.
+Hence, we have now met all preconditions of Theorem 1.7 with $x$ in place of $n$, $\varepsilon^{\prime}$ in place of $\varepsilon$, $p^{\prime}$ in place of $p$, and $H$ in place of $G$.
+It thus follows that whenever $k$ (and thus $n$ and hence $x$) are sufficiently large in terms of $\varepsilon$, the graph $H$ contains $F$ as a minor.
+In particular, the supergraph $G$ consisting of the red edges of the coloring of $K_{n}$ we initially considered contains $F$ as a minor.
+This is what we wanted to show, and concludes the proof in Case 1.
+Case 2. $H$ has vertex-connectivity less than $n(\log\log\log n)/(\log\log n)$.
+This means that there exists a subset of vertices $S\subseteq V(H)$ with $|S|<n(\log\log\log n)/(\log\log n)$ and a partition $(A,B)$ of $V(H)\setminus S$ into two non-empty parts such that there are no edges in $H$ connecting $A$ and $B$.
+Pick any vertices $a\in A$ and $b\in B$.
+Then all neighbors of $a$ lie in $A\cup S$, and all neighbors of $b$ lie in $B\cup S$.
+Since $H$ by definition has minimum degree at least $\varepsilon^{\prime}n$, it follows that $|A\cup S|,|B\cup S|\geq\varepsilon^{\prime}n$, and hence that
+$\min\{|A|,|B|\}\geq\varepsilon^{\prime}n-|S|>\left(\varepsilon^{\prime}-\frac{\log\log\log n}{\log\log n}\right)n\geq k$
+for $k$ sufficiently large.
+Observe that all edges of our initial coloring of $K_{n}$ spanned between $A$ and $B$ must be colored blue, since they cannot be contained in $G$.
+It follows that our initial coloring of $K_{n}$ contains a blue $K_{k,k}$ subgraph.
+Since $K_{k,k}$ contains $K_{k}$-minor, and hence also $F$-minor, we find that in Case 2 (provided $k$ is sufficiently large), the blue subgraph of our colored $K_{n}$ contains $F$ as a minor.
+This concludes the proof in Case 2.
+This concludes the proof of the desired upper bound on $R_{h}(F;2)$ and thus of the first part of the theorem.
 Now we prove the bounds for small $k$ and give general, but not optimal bounds that hold for any $k$.
-To see the lower bound of 2(k2)2(k-2), consider a union of two red vertex-disjoint cliques on k2k-2 vertices each so that all other edges are blue.
+To see the lower bound of 2k22k-2, consider a union of two red vertex-disjoint cliques, one on k1k-1 and the other on k2k-2 vertices, and color all other edges in between them blue.
 The upper bound follows from Theorem 1.16 by Bollobás [3], that claims that $c(k)\leq 8(k-2)\lfloor\log_{2}(k-2)\rfloor$.
-Indeed, let $G$ be a majority color class in a given red/blue coloring of the edges of $K_{n}$.
+Indeed, let $G$ be a majority color class in a given red/blue coloring of the edges of $K_{n}$ with $n=32(k-2)\lfloor\log(k-2)\rfloor+1$.
 It has at least n(n1)/4n(n-1)/4 edges.
@@
 Both $G$ and $\overline{G}$ are not bipartite since otherwise one of the parts has size at least $4$.
-By a special case of Hadwiger conjecture, $\chi(G)=\chi(\overline{G})=3$.
+By a special case of Hadwiger’s conjecture, $\chi(G)=\chi(\overline{G})=3$.
 Thus $G$ and $\overline{G}$ are three-partite graphs with parts of sizes at most three, i.e., of sizes $3,3,1$ or $3,2,2$.
@@
 Both the red and the blue graphs are three-chromatic with color classes of sizes $3,2,2$.
-This leaves a final not so long case analysis to demonstrate the existence of a monochromatic $K_{4}$-minor, a contradiction.
+This leaves a final, not too long, case analysis to demonstrate the existence of a monochromatic $K_{4}$-minor, a contradiction.
 ∎
@@
 3 Concluding remarks
-We showed that $R_{h}(k;2)=ck\sqrt{\log k}(1+o(1))$, where $c$ is a constant, $1\leq c\leq 1.03$....
-We believe that $c=1$.
-It would be interesting to find an easy direct proof of an upper bound on $R_{h}(k;2)\leq ck\sqrt{\log k}$, for some, maybe large constant $c$, but avoiding the density Theorem 1.1.
+We determined the asymptotic value of $R_{h}(k;\ell)$ when $\ell=2$ and when $k\gg\ell\gg 1$.
+It would be interesting to extend these results, and in particular asymptotically determine the value of $R_{h}(k;\ell)$ as $\ell\geq 3$ is fixed and $k\rightarrow\infty$.
+It would also be nice to find an easy direct proof of an upper bound on $R_{h}(k;2)\leq ck\sqrt{\log k}$, for some, maybe large constant $c$, but avoiding the density Theorem 1.1.
 To prove the lower bound on $R_{h}(k;\ell)$, one can also use the following packing result by taking $G$ being an $H$-factor, where $H$ is a graph on $k\sqrt{\log k}(1+o(1))$ vertices, density $\beta k\sqrt{\log k}$ and no $K_{k}$-minor.
-Pack $K_{n(1-\varepsilon)}$ with copies of $G$, give each such copy of $G$ own color.
-Decompose the remaining edges of $K_{n}$ into forests using Nash-Williams theorem and give each such forest a new own color.
+Pack $K_{n(1-\varepsilon)}$ with copies of $G$, give each such copy of $G$ its own color.
+Decompose the remaining edges of $K_{n}$ into forests using the Nash-Williams theorem and give each such forest a new own color.
 Theorem 3.1 (Messuti, Rödl, Schacht [10]).
@@
 If $G\in\mathcal{G}$, $\Delta(G)\leq\Delta$ and $|V(G)|\leq n$, and $\ell|E(G)|\leq\binom{n}{2}$ then $\ell$ copies of $G$ pack in $K_{(1+\varepsilon)n}$.
+Note that our Theorem 1.8 nails down the value of $R_{h}(F;2)$ precisely for every graph $F$ with $\gamma$ bounded away from zero, and thus for almost all graphs on $k$ vertices with at least $k^{1+\varepsilon}$ edges for any fixed $\varepsilon>0$ [11].
+However, in the sparse regime, i.e., when $F$ has only $k^{1+o(1)}$ edges, it only says that $R_{h}(F;2)=o(k\sqrt{\log k})$.
+Thus, it would be interesting to determine $R_{h}(F;2)$ more precisely for very sparse graphs, with, say, a constant average degree.
+A natural first step is to consider $R_{h}(F)$ for trees $F$ on $k$ vertices.
+In the case that $F=P_{k}$ is a path on $k$ vertices, we simply have $R_{h}(P_{k};2)=R(P_{k};2)=\left\lfloor\frac{3k-2}{2}\right\rfloor$ for every $k$, since any graph contains $P_{k}$ as a minor if and only if it contains $P_{k}$ as a subgraph, and since the Ramsey number of a path was determined to equal the above value by Gerencsér and Gyárfás [6].
+Moreover, also for multiple colors the known results on the Ramsey numbers of paths extend to their minor-Ramsey numbers.
+On the other extreme, for a star $S_{k}=K_{1,k}$ on $k+1$ vertices, we show that $R_{h}(S_{k};2)$ is by a factor two smaller than $R(S_{k};2)=2k+O(1)$.
+Proposition 3.2.
+We have $R_{h}(S_{k};2)=k+o(k)$.
+Since no graph on at most $k$ vertices contains $S_{k}$ a minor, we have $R_{h}(S_{k};2)\geq k+1$, and hence it suffices to prove the upper bound.
+So fix $\varepsilon>0$ arbitrarily and let us show that for every sufficiently large $k$, every red/blue-coloring of the edges of $K_{\lceil(1+\varepsilon)k\rceil}$ contains a monochromatic $S_{k}$-minor.
+In the following, we set $n:=\lceil(1+\varepsilon)k\rceil$.
+Consider any red/blue-coloring of the edges of $K_{n}$ and let $G_{r},G_{b}$ denote the spanning subgraphs of $K_{n}$ formed by the red and blue edges, respectively.
+If any vertex in any of $G_{r},G_{b}$ has degree at least $k$, then our coloring contains $S_{k}$ as a monohromatic subgraph and we are done.
+Hence, in the following we may assume that $\Delta(G_{r}),\Delta(G_{b})<k$.
+Since $G_{r}$ and $G_{b}$ are complements of each other, this in particular implies that $G_{r}$ and $G_{b}$ have minimum degree at least $(n-1)-(k-1)\geq\lceil\varepsilon k\rceil=:s$ each.
+Now, let $S\subseteq V(K_{n})$ be a random subset of vertices created as follows.
+Randomly pick vertices $v_{1},\ldots,v_{s}$ from $V(K_{n})$ with repetition, i.e.
+such that $v_{i}$ is uniformly random in $V(K_{n})$ independently from all other $v_{j}$.
+Finally, we set $S:=\{v_{1},\ldots,v_{s}\}$.
+We have $|S|\leq s$ and hence $|V(K_{n})\setminus S|\geq n-s=k$.
+We now claim that for every sufficiently large $k$, the set $S$ simultaneously forms a dominating set in both of the graphs $G_{r},G_{b}$.
+Indeed, for a fixed vertex $v\in V(K_{n})$ and a fixed color $i\in\{r,b\}$, the probability that $v$ has no neighbors in $S$ in the graph $G_{i}$, i.e., that none of the vertices $v_{1},\ldots,v_{s}$ was picked from $N_{G_{i}}(v)$, is
+$\left(1-\frac{d_{G_{i}}(v)}{n}\right)^{s}\leq\exp\left(-\frac{d_{G_{i}}(v)s}{n}\right)\leq\exp\left(-\frac{s^{2}}{n}\right)=\exp(-\Theta(k)).$
+Hence, the probability that some vertex $v$ is not dominated by $S$ in $G_{i}$ for some color $i\in\{r,b\}$ is at most $n\cdot 2\cdot\exp(-\Theta(k))=O(k)\cdot\exp(-\Theta(k))\rightarrow 0$ for $k\rightarrow\infty$.
+Hence, indeed for every sufficiently large $k$, with positive probability, the set $S$ is a dominating set of both $G_{r}$ and $G_{b}$.
+Now, assuming $k$ is sufficiently large, fix a set $S$ of size at most $s$ which dominates $G_{r}$ and $G_{b}$.
+Since $G_{r}[S]$ and $G_{b}[S]$ are complementary graphs on vertex set $S$, at least one of them, say $G_{r}[S]$, must be connected.
+But then we obtain $S_{k}$ as a minor of $G_{r}$ by contracting $S$ (more precisely, all edges contained in it) into a single vertex.
+This is the desired outcome and concludes the proof.
+∎
 4 Acknowledgements
-The author is grateful to Noga Alon for several very fruitful discussions and suggestions. She also thanks Jan Malte Weirich for discussions.
+The first author is grateful to Noga Alon for several very fruitful discussions and suggestions. She also thanks Jan Malte Weirich for discussions.
+The second author gratefully acknowledges funding from the Ambizione grant No. 216071 of the Swiss National Science Foundation.